OpenStudy (anonymous):

solve 3√5-2√7+√45-√28 i got 2√12 but i think I am completely wrong?

OpenStudy (anonymous):

\[\sqrt{45}=3\sqrt{5}\] \[\sqrt{28}=2\sqrt{7} \]

OpenStudy (anonymous):

okay let me redo the problem